OpenStudy (anonymous):

Mr. Doodle’s grade distribution over the past 3 years for a course in college algebra is shown in the chart below. Grade Number A 45 B 180 C 210 D 95 F 65 I 5 If Jane plans to take a college algebra course with Mr. Doodle, determine the empirical probability that she receives a C.

OpenStudy (anonymous):

empirical probability would be number of C received divided by the total number of grades given

OpenStudy (anonymous):

this what I started with:45+180+210+5=65+5=500 c=110 100/500 am I on the right track?????????

OpenStudy (anonymous):

I mean 110/500

OpenStudy (anonymous):

C has 110 or 210??

OpenStudy (anonymous):

ok but am I doing it correctly? if no can you explain what I'm doing wrong?

OpenStudy (anonymous):

Yes you are doing right...

OpenStudy (kropot72):

The total number of grades given over the past 3 years is 600

OpenStudy (anonymous):

ok I really appreciate the help

OpenStudy (anonymous):

@kropot good eye....thanks